OpenStudy (anonymous):

Easy question, i just suck at math. HELP PLEASE! Sam has 15 coins, all nickels and dimes, with a total value of $0.90. How many of each coin does he have? (Let d = the number of dimes and n = the number of nickels). A. Write a system of equations to model the problem. B. Solve the system. C. State your answer with the proper label or labels.

OpenStudy (johnweldon1993):

It is a system of equations 15 coins in all...all nickels and dimes,,,let d = dimes and n = nickels d + n = 15 The total amount of money = .90 cents so .10d + .05n = .90 our 2 equations d + n = 15 .10d + .05n = .90 can you solve it from here?

OpenStudy (anonymous):

Can you please actually solve it for me? I really am horrible at math, and I can't do this one at all.

OpenStudy (texaschic101):

d + n = 15 --> d = 15 - n .10d + .05n = .90 sub 15 - n in for d in the 2nd equation .10d + .05n = .90 .10(15 - n) + .05n = .90 (distribute through the parenthesis) 1.50 - .10n + .05n = .90 (combine like terms) 1.50 - .05n = .90 -.05n = .90 - 1.50 -.05n = - .60 n = - .60/-.05 n = 12 now sub 12 into the first equation n + d = 15 12 + d = 15 d = 15 - 12 d = 3 check... .05n + .10d = .90 .05(12) + .10(3) = .90 .60 + .30 = .90 .90 = .90 (correct) There is 3 dimes and 12 nickels

OpenStudy (johnweldon1993):

Excellent work @texaschic101 :)

OpenStudy (texaschic101):

thank you....I probably should not have given the answer....I was just giving @PuppyPawsB a little break :)

OpenStudy (anonymous):

thanks so much @texaschick101